angle 1 and angle 2 are supplementary angles. m of angle 1= x-39 and measure of angle 2= x+61. Find the measure of each angle.
Answer by cleomenius(959)

x - 39 + x + 61 = 180
2x + 22 = 180
2x = 158
x = 79
79 - 39 = 40 degrees.
79 + 61 = 140 degrees.
40 + 140 - 180 degrees.
This does check.
Cleomenius.